Find the solution set for the equation. Please help?!!
|7w+3|-12>=-12

Answer by jim_thompson5910(35256)

|7w+3|-12>=-12

|7w+3|>=-12+12

|7w+3| >= 0

7w+3 >= 0 or 7w+3 <= 0

7w >= -3 or 7w <= -3

w >= -3/7 or w <= -3/7

The two inequalities combine to form the entire real number line, so that's why the solution set is the set of all real numbers.